# Transfer Learning for High-dimensional Quantile Regression with Distribution Shift

### v1

@@ 5.3 Inference Results
Table 1: Coverage probability of $95\%$ confidence intervals of various methods with $\delta^{(k)}_{1}=0.01$
@@
Debias_tar 0.953 0.939 0.951
Debias_trans_pool 0.917 0.934 0.909
Debias_trans_tar 0.955 0.950 0.962
Debias_trans 0.951 0.954 0.945
Debias_dl 0.938 0.954 0.961
Table 2: Coverage probability of $95\%$ confidence intervals of various methods with $\delta^{(k)}_{1}=0.1$
@@
Debias_tar 0.953 0.939 0.951
Debias_trans_pool 0.789 0.766 0.756
Debias_trans_tar 0.96 0.96 0.96
Debias_trans 0.846 0.809 0.817
Debias_dl 0.951 0.962 0.961
Figure 5.3: Inference results of various methods with $\delta^{(k)}_{1}=0.01$
Figure 5.4: Inference results of various methods with $\delta^{(k)}_{1}=0.1$
From Table 1-2 and Figure 5.3-5.4, we highlight three points.
  1. 1.
    From Table 1, our method Debias_trans enjoys the better coverage performance, which is the nearest to $0.95$ among all methods. Since Debias_tar, Debias_trans_tar and Debias_dl only solve the target score functions approach, their coverage probabilities are somewhat fluctuant around $0.95$. The Debias_trans_pool approach has the worst coverage properties, which again demonstrates that introducing source stuides with large residual shift will bring negative effect to the transfer efficiency.
  2. 2.
    From the estimation bias and interval length in Figure 5.3, we can see that Debias_trans have less bias and shorter interval lengths than Debias_trans_tar and Debias_tar, collaborating to the theory that our approach has faster asymptotic rates by incorporating informative sources. Although Debias_trans_pool also has the comparable biases and intervals (utilizing the most studies), it sacrificed coverage accuracy due to residual shift. Moreover, since Debias_dl debiasing on the target study, it only obtains $\sqrt{n_{0}}$-rate and hence has relatively larger bias and longer interval lengths.
  3. 3.
    By comparing the density plots with the standard normal (orange dashed line) in Figure 5.3, it can be seen that normalized estimates of all these methods are close to the standard normal, except that Debias_trans_pool at $\tau=0.7$, Debias_tar, Debias_trans_tar at $\tau=0.2$ lack a little bit efficiency. Therefore, under small $\delta_{1}^{(k)}$ our transfer approach can explore source information validly and conduct more efficient inference than previous works, which will offer more insights for statistical inference under knowledge transfer.
  4. 4.
    From Table 2 and Figure 5.4, when $|\delta_{1}^{(k)}|$ is large, unfortunately our Debias_trans performs much worse than the target debiasing methods, which is in our expectation since Condition 8 is no longer satisfied. The density plots in Figure 5.4 show that the transfer debiased estimate is no longer normal, and the Shapiro-Wilk test rejects the normality hypothesis. In a similar way, the bootstrap approach in Algorithm 4 can also indicate us that debiasing on source studies will generate unreliable estimates. Therefore, we recommend only debiasing on the target study, namely the Debias_trans_tar approach, which can keep the coverage proportions with more stable interval lengths than Debias_dl.
6 Application
The proposed transfer learning algorithm is applied to the Genotype-Tissue Expression (GTEx) data, available at https://gtexportal.org/. This dataset contains gene expression levels from 49 tissues, involving 838 individuals, and includes 1,207,976 observations for 38,187 genes. Following the previous work (Li et al.,, 2022; Zhang and Zhu,, 2023; Zhang et al.,, 2024), our study examines gene regulation within the central nervous system (CNS) across various tissues. The CNS-related genes are grouped under MODULE_137, which includes 545 genes, along with an additional 1,632 genes that show significant enrichment in the same experiments as those in the module. Please refer to https://www.gsea-msigdb.org/gsea/msigdb/cards/MODULE_137.html for a detailed description of this module.
Our focus is on the expression levels of the genes JAM2 and SH2D2A at the lower quantile. Specifically, we aim to predict the expression levels of these two genes at $\tau=0.2$ within a target tissue by leveraging the expression of other central nervous system (CNS) genes. Recent research has identified that, JAM2 and SH2D2A play a significant role in various cellular functions and disease progressions, particularly in cancer and immune response.
@@ 6 Application
  • JAM2 (Junctional Adhesion Molecule 2) is a gene that encodes a protein involved in tight junctions between epithelial and endothelial cells, playing a crucial role in cell adhesion and maintaining barrier functions. It is associated with various cellular processes such as cell proliferation and migration. Studies have shown that low expression of JAM2 is linked to poor prognosis in several cancers, including lung adenocarcinoma (LUAD), where reduced JAM2 levels correlate with tumor progression and metastasis. Investigating its low expression helps in understanding cancer progression and potentially identifying therapeutic targets for limiting tumor growth (Dong et al.,, 2024).
@@
  • SH2D2A encodes the T-cell-specific adapter protein (TSAd), which is essential for T-cell signaling and activation. The expression of SH2D2A is tightly regulated at both the transcriptional and translational levels, with cAMP signaling shown to induce its mRNA expression in T cells. However, low levels of SH2D2A protein may affect immune responses. Notably, reduced SH2D2A expression can enhance T-cell-mediated anti-tumor immunity, making it an interesting target in cancer immunotherapy research (Berge et al.,, 2012).
We analyze 13 brain tissues as separate target studies, estimating the model for each tissue individually. The target data is divided into five folds, where each fold is used as a prediction set, while the remaining four folds are used for training. The methods L1-QR, Pooling, TransQR_2step, and TransQR from the simulation section are applied in this analysis. Detailed data pre-processing and tuning parameter selection can be found in the supplementary materials. Figure 6.1 shows the relative prediction errors of Pooling, TransQR_2step, and our approach, TransQR, compared to L1-QR across the target tissues, with the black horizontal line representing a ratio of 1. As expected, TransQR consistently achieves the lowest prediction errors, significantly improving the performance of the target task in nearly all cases. We observe that the two-step approach, TransQR_2step, tends to select more source studies than TransQR, indicating that our transferable set detection can effectively identify valuable studies even under parameter or residual shift. Moreover, in some instances, TransQR_2step does not provide substantial improvements over Pooling and L1-QR, further demonstrating the robustness of our transfer framework in addressing covariate shift. Compared to the mean regression results reported in Li et al., (2022); Zhao et al., (2023), our results offer richer insights for understanding the whole distribution of gene expression.
Figure 6.1: Relative prediction errors for the expression level of gene JAM2 and SH2D2A under quantile level $\tau=0.2$.
7 Discussion
To tackle the possible negative transfer caused by distribution shift, we introduce a novel transferable set with screening both $\|\boldsymbol{\delta}^{(k)}\|_{1}$ and $\mathbb{E}[f_{i}^{(k)}]$ to address the parameter and residual shift, also we adopt the $\ell_{1}$-minimization framework to avoid the trouble from covariate shift. Along with the transferable set detection and parameter estimation, we propose an orthogonal debiased approach for statistical inference with leveraging informative sources. Theoretical results as well as numerical experiments validate the effectiveness of our proposed method. There are several avenues to generalize our method, such as
  • We can construct source weights proportional to their respective $\tau$-th densities to achieve sharper results. Leveraging the density at objective quantiles to enhance the efficiency of estimators has been a common approach in the literature on weighted quantile regression (Jiang et al.,, 2012; Huang et al.,, 2015; Lu and Fan,, 2015; Xiong and Tian,, 2022), as the Hessian matrices of check loss functions are directly proportional to $\tau$-th densities. Furthermore, abundant source information at multiple quantiles can be utilized to predict the target study’s objective quantile coefficient. For instance, we can extend panel quantile regression (Zhang et al.,, 2019) or GMM quantile regression (Firpo et al.,, 2022) to the transfer setting.
  • Our transferable set adopts a common all-in-or-all-out manner, thus in (3) we use the expectation of $\tau$-th densities as the transferability measure. In fact, we can generalize this measure to select partial informative data from each source study, especially when the density is strongly dependent on the covariates. To be specific, define the transferable dataset (not transferable set) $\mathcal{C}^{\prime}_{h}$ as
    $\mathcal{C}^{\prime}_{h}=\left\{(\boldsymbol{x}_{i}^{(k)},y_{i}^{(k)}):\quad 1\leq k\leq K,\quad\|\boldsymbol{\delta}^{(k)}\|_{1}\leq h_{1},\quad\frac{1}{f^{(k)}_{i}}\leq h_{2}\frac{1}{f^{(0)}_{i}}\right\}.$
    Then we use the data in $\mathcal{C}^{\prime}_{h}$ to conduct knowledge transfer, where we can leverage more information from related source studies.
  • Due to the privacy or data storage concern, we may not be able to combine the individual-level data of all studies in a single machine. Assume we can only transmit summary statistics between studies, and the final result is calculated on the machine with target data (referred as the target machine). By carefully observing our Algorithm 2, we can see that almost every step is based on the local data and the summary statistics from other studies. Briefly speaking, we just need to transmit the estimators like $\widetilde{\boldsymbol{\beta}},\{\widetilde{\boldsymbol{\delta}}^{(k)}\}_{k=1}^{K}$ between each source machine and the target machine. The target machine can then determine the transferable set $\widehat{\mathcal{C}}$. The only step needed to pool the data from different studies is the last step of Algorithm 2. To accommodate this step to the distributed setting, we can use existing algorithms (for example Algorithm 1 in Chen et al., (2020)) to conduct distributed high-dimensional quantile regression. Therefore, our algorithm can be easily generalized to the communication-efficient version with multiple data storage nodes.
There are many interesting topics of transfer learning for future research. For example, distributionally robust transfer learning aims to obtain an optimized target distribution in a prespecified set with feasible distributions related to the source distributions. The present works mainly focus on the feasible distribution set constructed by the convex hull of source distributions (Xiong et al.,, 2023; Wang et al.,, 2023). We can extend the feasible set in terms of the Wasserstein distance metric (Blanchet et al.,, 2019; Gao and Kleywegt,, 2023), which may help us to select source domains more effectively, with learning a target model more robust to various distribution shifts.
Acknowledgment
The research of Zhu is partially supported by the National Natural Science Foundation of China 12071087, 12331009.

### v2

@@ 8.2 Additional Simulation Results
Table 2: Coverage probability of $95\%$ confidence intervals (CP), average absolute bias (Bias), standard error (SE), and average estimated standard error (ESE) of various methods with $|\delta^{(k)}_{1}|=0.1$.
@@
CP Bias SE ESE CP Bias SE ESE CP Bias SE ESE
Debias_tar 0.965 0.101 0.123 0.138 0.954 0.0857 0.108 0.110 0.953 0.0880 0.110 0.117
Debias_trans_pool 0.809 0.0892 0.107 0.0738 0.788 0.0792 0.0955 0.0627 0.781 0.0823 0.0998 0.0657
Debias_trans_tar 0.971 0.0958 0.125 0.138 0.950 0.0844 0.105 0.110 0.954 0.0875 0.110 0.117
Debias_trans 0.866 0.0861 0.104 0.0807 0.836 0.0772 0.0937 0.0677 0.853 0.0791 0.0953 0.0711
Debias_dl 0.928 0.0876 0.109 0.104 0.935 0.0714 0.0893 0.0854 0.936 0.0782 0.0964 0.0910
Figure 15: Inference results of various methods with $|\delta^{(k)}_{1}|=0.1$, including boxplots of bias, boxplots of confidence interval lengths, and density plots of normalized estimates compared to $\mathcal{N}(0,1)$ with black dotted line.
Table 3: Coverage probability of $95\%$ confidence intervals (CP), average absolute bias (Bias), standard error (SE), and average estimated standard error (ESE) of various methods with strong covariate shift.
@@
CP Bias SE ESE CP Bias SE ESE CP Bias SE ESE
Debias_tar 0.963 0.103 0.137 0.138 0.941 0.0853 0.108 0.110 0.954 0.0934 0.114 0.118
Debias_trans_pool 0.927 0.0620 0.0778 0.0696 0.944 0.0556 0.0687 0.0656 0.918 0.0594 0.0774 0.0655
Debias_trans_tar 0.979 0.0964 0.134 0.138 0.952 0.0826 0.105 0.110 0.963 0.0879 0.110 0.118
Debias_trans 0.961 0.0592 0.0755 0.0769 0.963 0.0515 0.0652 0.0710 0.948 0.0565 0.0784 0.0713
Debias_dl 0.966 0.0784 0.0986 0.110 0.945 0.0677 0.0868 0.0884 0.951 0.0724 0.0914 0.0962
Figure 16: Inference results of various methods with strong covariate shift, including boxplots of bias, boxplots of confidence interval lengths, and density plots of normalized estimates compared to $\mathcal{N}(0,1)$ with black dotted line.
Table 4: Coverage probability of $95\%$ confidence intervals (CP), average absolute bias (Bias), standard error (SE), and average estimated standard error (ESE) of various methods under $t$-distribution with strong covariate shift.
Method $\tau=0.2$ $\tau=0.5$ $\tau=0.7$
CP Bias SE ESE CP Bias SE ESE CP Bias SE ESE
Debias_tar 0.929 0.140 0.166 0.165 0.957 0.100 0.124 0.131 0.926 0.116 0.162 0.137
Debias_trans_pool 0.905 0.0795 0.107 0.0833 0.933 0.0624 0.0802 0.0801 0.906 0.0689 0.0893 0.0770
Debias_trans_tar 0.945 0.128 0.161 0.165 0.966 0.0943 0.119 0.131 0.946 0.107 0.135 0.137
Debias_trans 0.944 0.0753 0.105 0.0935 0.978 0.0588 0.0745 0.0874 0.948 0.0648 0.0836 0.0844
Debias_dl 0.940 0.110 0.135 0.138 0.956 0.0787 0.0996 0.101 0.952 0.0857 0.108 0.113
Figure 17: Inference results of various methods under $t$-distribution with strong covariate shift, including boxplots of bias, boxplots of confidence interval lengths, and density plots of normalized estimates compared to $\mathcal{N}(0,1)$ with black dotted line.
8.3 Details of Real Data Application
We first give a brief introduction for the two genes studied in Section 6.
@@ 8.3 Details of Real Data Application
  • JAM2 (Junctional Adhesion Molecule 2) is a gene that encodes a protein involved in tight junctions between epithelial and endothelial cells, playing a crucial role in cell adhesion and maintaining barrier functions. It is associated with various cellular processes such as cell proliferation and migration. Studies have shown that low expression of JAM2 is linked to poor prognosis in several cancers, including lung adenocarcinoma (LUAD), where reduced JAM2 levels correlate with tumor progression and metastasis. Investigating its low expression helps in understanding cancer progression and identifying therapeutic targets for limiting tumor growth [dong2024jam2].
@@
  • SH2D2A encodes the T-cell-specific adapter protein (TSAd), which is essential for T-cell signaling and activation. The expression of SH2D2A is tightly regulated at both the transcriptional and translational levels, with cAMP signaling shown to induce its mRNA expression in T cells. However, low levels of SH2D2A protein may affect immune responses. Notably, reduced SH2D2A expression can enhance T-cell-mediated anti-tumor immunity, making it an interesting target in cancer immunotherapy research [berge2012sh2d2a].
Then we provide the complete name for each target issue, which is shown in Table 5.
Table 5: Complete Name for 11 Target Brain Issues in the Application
Tissue Abbreviation
brain amygdala Amygdala
brain hypothalamus Hypothalamus
brain caudate basal ganglia C.B.ganglia
brain cerebellar hemisphere C.hemisphere
brain cerebellum Cerebellum
brain cortex Cortex
brain frontal cortex ba9 F.cortex
brain anterior cingulate cortex ba24 A.C.cortex
brain nucleus accumbens basal ganglia N.C.A.B.ganglia
brain putamen basal ganglia P.B.ganglia
brain spinal cord cervical c-1 S.C.cervical
There are some additional points in our practical implementation:
  • The gene expression data is pre-processed by filtering out genes with constant expression levels and applying standard normalization.
  • We removed the source tissues with sample sizes less than $100$.
  • We performed threshold value searching for the “brain cortex” tissue as the target study in real data applications. The five-fold cross-validation relative prediction errors on the “brain cortex” tissue, shown in Figure 18, guide our choice of $t_{1}=7$ and $t_{2}=0.8$ for all tasks. The other regularization parameters follow the guidance in Section 8.1.3.
  • We observe that source data appears less beneficial at $\tau=0.5$ but seems to be quite helpful at $\tau=0.2$. This suggests that the distribution shift between the target and source studies may vary with the quantile level of model responses.
Figure 18: Transfer performance under different values for the thresholds $t_{1},t_{2}$ for predicting the cortex tissue at quantile level $\tau=0.2$.